OpenStudy (candy13106):

-15=6p+15-10p solve the equation

OpenStudy (cj49):

p=15/2

OpenStudy (candy13106):

how did u get that

OpenStudy (cj49):

-15=4p+15 4p-30=0 4p=30 p=30/4 p=15/2

OpenStudy (candy13106):

idk i got 7.5

OpenStudy (candy13106):

@thomaster

OpenStudy (cj49):

can u show wat u did

OpenStudy (candy13106):

idk how i got it

OpenStudy (jessiegonzales):

its 7.5 15/2

OpenStudy (candy13106):

its 7.5 but how?

OpenStudy (jessiegonzales):

didn't someone just help you on this? Anyway here it is again.. -15=6p+15-10p -30=6p-10p -30=-4p -30/-4=p p=7.5